OpenStudy (anonymous):

The probability of Anna, Bianca and Cilem passing a geography test is 0.25, 0.5 and 0.4 respectively. The probability that only two girls will past the next geography test is?

OpenStudy (chaise):

If Cilem is a boy. 0.25*0.5=1/8 or 0.125

OpenStudy (anonymous):

P (A and B) = P(a) * P(B) = 0.25 * 0.5 = 0.125

OpenStudy (anonymous):

the answer is 0.275

OpenStudy (anonymous):

Assuming Cilem is also a girl... then we have a tricky problem that only 2 out of the three pass the test, which means we need to examine all permutations. Call the three A, B and C we have three permutations: A passes, B passes, C fails A passes, C passes, B fails B passes, C passes, A fails We are given the chances of passing. The chances of failing a just 1 - chance of passing. Using ! for not passing we get: AxBx!C = .25 x .5 x .6 = .075 AxCx!B = .25 x .4 x .5 = .05 BxCx!A = .5 x .4 x .75 = .15 Adding we get .275

OpenStudy (anonymous):

oh i see thank you very much

OpenStudy (anonymous):

Thanks fo the medal!

OpenStudy (anonymous):

nw:)